A point charge Q1=3uc at origin and a point charge Q2=-7 uc is placed at x= 0.4m on the x axis or cartesian system.determine the electric field E(x)=E(x)x at all points along x axis
1
Expert's answer
2020-05-18T10:13:44-0400

Let's draw the rough picture of the charges lies on the X axis,


Clearly, electric field due to charge "Q_1=3 \mu C" at "R" is


"E_1=\frac{1}{4 \pi \epsilon_0}\frac{Q_1}{(x+0.4)^2} \hspace{1cm}(1)"

similarly, electric field due to charge "Q_2=-7 \mu C" at "R" is

"E_2= \frac{1}{4 \pi \epsilon_0}\frac{Q_2}{x^2} \hspace{1cm}(2)"

Now, by principle of superposition total electric field at "R" will be


"E_{total}=\frac{1}{4 \pi \epsilon_0}\bigg(\frac{Q_1}{(x+0.4)^2}+\frac{Q_2}{x^2} \bigg)\\ \implies
E_{total}=\frac{1}{4 \pi \epsilon_0}\bigg(\frac{3\times 10^{-6}}{(x+0.4)^2}+\frac{-7 \times 10^{-6}}{x^2} \bigg)N/C\\ 
\implies E_{total}=\frac{1\times10^{-6}}{4 \pi \epsilon_0}\bigg(\frac{3}{(x+0.4)^2}-\frac{7 }{x^2} \bigg) N/C\\"

Direction of "E_{total}" is along the negative "x" axis.


Hence,we are done.